# Adaptive Rank, Reduced Forgetting: Knowledge Retention in Continual Learning Vision-Language Models with Dynamic Rank-Selective LoRA

### v5

@@ 3.3 Proposed CoDyRA
   Zero-shot [42] 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
   LwF [29] 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
   LwF-VR [13] 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
   WiSE-FT [54] 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
   ZSCL [66] 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
@@
   LwF [29] 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
   LwF-VR [13] 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
   WiSE-FT [54] 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
   ZSCL [66] 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
@@
   LwF [29] 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
   LwF-VR [13] 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
   WiSE-FT [54] 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
   ZSCL [66] 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
@@
4 Experiments
4.1 Experimental Setting
Datasets. We evaluate our method on two widely used continual learning settings for vision-language models: MTIL [66, 61] and X-TAIL [58]. For the MTIL setting, a total of 11 datasets are used, where each dataset is treated as an individual task to be incrementally learned. For the X-TAIL setting, CIFAR100 is excluded to prevent domain overlap, following the protocol in [58]. In line with [58], we use a 5-shot split for MTIL and a 16-shot split for X-TAIL.
Evaluation Metrics. We adopt the same evaluation metrics as prior works, including “Transfer”, “Average” , and “Last”. “Transfer”evaluates the model’s zero-shot transfer capability on unseen data. “Last” assesses the model’s ability to retain knowledge from earlier tasks, and“Average” provides a composite score, representing the mean performance across both the “Transfer” and “Last”.
Implementation Details. We follow the experimental setups in [66, 61, 58] and use the CLIP model with a ViT-B/16 backbone [42] for all experiments. By default, CoDyRA is applied to all pre-trained weight matrices in both the vision and text encoders, with an initial rank of 16. Each task in both experimental settings is trained over 500 iterations using the AdamW optimizer [33]. The dense training iteration ratio is set to 0.5 for X-TAIL and 0.7 for MTIL, relative to the total training iterations. The maximum threshold for adaptive pruning $\kappa$ is fixed at 0.005 for both settings.
4.2 Experimental Results
Cross-domain task-agnostic incremental learning. We report the X-TAIL setting evaluation results in Table 1. In contrast to prior methods [61, 58] , which rely on zero-shot predictions to preserve pre-trained capabilities and mitigate forgetting, our method overcomes this limitation by continually refining the model. Notably, while these approaches are bounded by the upper limit of zero-shot performance, our approach surpasses this limit, achieving a higher average transfer accuracy. Although some predictions on unseen data might be slightly affected due to the continual parameter updates, our method ultimately improves overall transfer accuracy by adapting the pre-trained model to new tasks while enhancing its generalization capabilities.
[…]
Multi-domain task-incremental learning. We further evaluate our proposed method under the few-shot MTIL setting, as shown in Table 2, following [61, 58]. Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this scenario. Without incorporating any additional techniques, our approach outperforms previous methods across all evaluation metrics, achieving an average transfer accuracy that surpasses the upper bound set by zero-shot predictions in prior works. Furthermore, by integrating techniques from previous approaches [61, 58], our method can be further enhanced, leading to even greater performance gains.
@@ 4.3 Discussions
Pre-trained knowledge retention and potential pre-trained model improvement. To understand how our method improves transfer performance beyond the upper bound set by zero-shot predictions in prior works, we evaluated on hold-out datasets never encountered during continual learning, specifically CIFAR100 [27], Places365 [67], and ImageNet-1k [10], as shown in Table 3.
Method Average Transfer Accuracy on Unseen Data
CIFAR100 Places365 ImageNet-1k Average
Zero-Shot [42] 68.24 33.77 66.72 56.24
MoE-Adapter [61] 68.24 33.77 66.72 56.24
RAIL [58] 68.24 33.77 66.72 56.24
CoDyRA 68.95 36.52 68.52 58.00
Unlike prior approaches [61, 58], which rely exclusively on zero-shot predictions for unseen data—thereby constraining their performance to that of the original CLIP model—our method overcomes this limitation. By effectively integrating new knowledge into the pre-trained parameter space while preserving existing capabilities, our approach enhances overall model performance. As a result, it achieves superior zero-shot transfer accuracy, surpassing the zero-shot performance limits of prior works. This demonstrates that CoDyRA effectively incorporates general knowledge into its knowledge base, further improving its overall representation capability.
@@
Rank allocation across different datasets. We visualize and analyze the distribution of activated ranks merged into the pre-trained weights across different datasets in Fig. 6 and Fig. 6. Overall, the Aircraft dataset generally requires more rank updates in the PTM than the Pets dataset.
@@ Methods
LWF [29] 129.6M None
ZSCL [66] 129.6M None
MoE-Adapters [61] 59.8M 13.35M
RAIL [58] N/A 24.18M / 9.01M
@@
In contrast, for the Pets dataset, more ranks are allocated to the Attention module’s Value and Output projections and the MLP’s Projection layer within the vision encoder (Fig. 6(b)). Additionally, rank allocation is primarily concentrated in the first and last layers (Fig. 6(c)), indicating a distinct adaptation pattern that balances knowledge retention and adaptation across datasets. These observations highlight the adaptive nature of our approach, where rank allocation dynamically adjusts based on dataset characteristics, ensuring effective domain adaptation while preserving pre-trained knowledge. We provide more visualizations and analyses in Appendix.
Computation cost. We compare the computational costs in terms of trainable parameters and additional inference requirements, as summarized in Fig. 4. For MoE-Adapter, the difference between training and testing parameters arises from its mixture-of-experts design: during training, all experts are active, whereas during inference, only the top-2 experts are selected, reducing the number of active parameters. RAIL-Primal incurs additional inference overhead due to its use of high-dimensional feature expansion through projection and regression layers. Meanwhile, RAIL-DUAL significantly increases memory usage by maintaining a memory bank of image features for all training samples across 1,100 classes. Notably, these methods experience linearly growing computational costs as the number of tasks increases. In contrast, our method minimizes training overhead by selectively retaining only ranks with high importance. Once training is complete, these updates are merged into the pre-trained weights, ensuring zero additional memory overhead during inference.
@@ 4.4 Ablation Studies
Effects of insertion locations and number of initial ranks. We examine the impact of different insertion locations and initial ranks in our method, with results presented in Fig. 7. These findings are consistent with our analyses shown in Fig. 4. Our results indicate that when training is restricted to the vision encoder, transfer accuracy drops significantly, suggesting that the pre-trained capabilities are negatively impacted. Moreover, at specific locations, increasing the maximum rank improves adaptation ability while not always reducing transfer accuracy.
Interestingly, restricting updates to only the Attention modules can surpass our reported state-of-the-art Transfer metric in Table 1 and Table 2. However, this comes at the expense of reduced adaptation capabilities. To strike a better balance between adaptation and retention, we enable updates across all modules by default, ensuring an optimal trade-off between learning new tasks and retaining the model’s pre-trained capabilities.
Analyses of pruning threshold and dense-iteration ratio. To better understand the trade-off between downstream learning and forgetting, we control the number of adaptive ranks by varying the maximum threshold $\kappa_{\max}$ and the fraction of dense training iterations (Fig. 8). Gradually raising $\kappa_{\max}$ improves Last accuracy by pruning more useless or noisy ranks, whereas overly aggressive pruning limits task adaptation; a modest setting of 0.005 strikes a good balance. The dense-iteration ratio governs pruning onset: pruning from the start (ratio = 0) removes ranks before it can learn from the training data, while delaying pruning too long increases the risk of retaining useless or noisy ranks that do not get pruned eventually.
@@ 5 Conclusion
In this paper, we introduced CoDyRA, a novel approach to addressing continual learning challenges in vision-language models. By dynamically adjusting ranks based on their importance, our method optimizes knowledge retention while enabling efficient adaptation to new tasks without adding inference overhead. Extensive evaluations across various continual learning settings demonstrate that our method achieves state-of-the-art performance, effectively balancing downstream adaptation and pre-trained retention.
Limitations and future works. Our approach merges updates into pre-trained weights without modeling inter-task correlations, mainly evaluated using pre-trained CLIP model. Future works could explore leveraging inter-task relationships and extending our method to other PTMs and settings.
Acknowledgments and Disclosure of Funding
This work was partially supported by the ARC DECRA Fellowship (DE230101591) awarded to D. Gong. H. Lu is affiliated with CSIRO Data61 through a PhD scholarship and acknowledges the support of the Google PhD Fellowship.
@@ Appendix A Additional Details of the Proposed Method
We apply proximal gradient descent [4] to handle the non-differentiable $\ell_{1}$ regularization of the importance weights [12]. At the $t$-th iteration, the update rule for the importance weights of each LoRA becomes:
@@
such that the input to the operator $\mathcal{T}_{\kappa}$ is the elements in $\hat{\mathbf{w}}^{m}_{t}$, $\hat{\mathbf{w}}^{m}_{t}$ is the value of $\mathbf{w}^{m}_{t}$ after the gradient update using only $\mathcal{L}_{\text{sup}}(\Delta_{t})$, and $\kappa=\eta_{t}\cdot\lambda$ is the threshold. The indicator function $\mathbbm{1}(\cdot)$ returns 1 if the condition is met, and 0 otherwise, while $\mathrm{sign}(\cdot)$ denotes the sign (+/)(+/-)( + / - ) of the input. This concludes our derivation of the dynamic update for the importance weights as presented in the main paper Eq. (4).
@@ Appendix B More Experimental Results and Ablations
Reproducibility Statement. The source code will be made publicly available upon acceptance of the paper. We report experiment results averaged over 3 independent runs for reproducibility. Experiments were conducted on a single NVIDIA GeForce RTX 3090 GPU.
B.1 Effects of Additional Adaptive Prediction Head on Visual Representations

### v6

@@ 4.2 Experimental Results
   Zero-shot (Radford et al., 2021) 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
   LwF (Li & Hoiem, 2017) 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
   LwF-VR (Ding et al., 2022) 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
   WiSE-FT (Wortsman et al., 2022) 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
   ZSCL (Zheng et al., 2023) 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
@@
   LwF (Li & Hoiem, 2017) 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
   LwF-VR (Ding et al., 2022) 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
   WiSE-FT (Wortsman et al., 2022) 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
   ZSCL (Zheng et al., 2023) 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
@@
   LwF (Li & Hoiem, 2017) 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
   LwF-VR (Ding et al., 2022) 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
   WiSE-FT (Wortsman et al., 2022) 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
   ZSCL (Zheng et al., 2023) 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
@@
Multi-domain task-incremental learning. We further evaluate our method in the few-shot MTIL setting (Table 3), following (Yu et al., 2024; Xu et al., 2024). As with X-TAIL, our approach outperforms prior methods across all metrics without extra techniques and even surpasses their zero-shot upper bound in average transfer accuracy. Incorporating techniques from (Yu et al., 2024; Xu et al., 2024) further boosts performance.
@@ 4.3 Discussions
Pre-trained knowledge retention and potential pre-trained model improvement. To assess transfer beyond the zero-shot limits of prior works, we evaluated on unseen datasets—CIFAR100 (Krizhevsky et al., 2009), Places365 (Zhou et al., 2017), and ImageNet-1k (Deng et al., 2009) (Table 2). Unlike methods that rely solely on zero-shot predictions (Yu et al., 2024; Xu et al., 2024) and are constrained by CLIP’s baseline, our approach integrates new knowledge into the pre-trained space while preserving existing capabilities. This yields superior zero-shot transfer accuracy, demonstrating that CoDyRA effectively expands general knowledge and improves overall representations.
Rank allocation across different datasets. We visualize the ranks merged into the pre-trained weights for different datasets in Fig. 6 and Fig. 6 (more in Appendix Fig. 11 to Fig. 17). The results show that CoDyRA adapts the number of active ranks differently across locations and tasks, reflecting the varying demands of each dataset. These strike a balance between plasticity and stability, demonstrating the method’s ability to regulate forgetting while enabling effective task learning.
In contrast, for the Pets dataset, more ranks are allocated to the Attention module’s Value and Output projections and the MLP’s Projection layer within the vision encoder (Fig. 6(b)). Additionally, rank allocation is primarily concentrated in the first and last layers (Fig. 6(c)), indicating a distinct adaptation pattern that balances knowledge retention and adaptation across datasets. These observations highlight the adaptive nature of our approach, where rank allocation dynamically adjusts based on dataset characteristics, ensuring effective domain adaptation while preserving pre-trained knowledge. We provide more visualizations and analyses in Appendix B.4.
@@ Methods
LWF (Li & Hoiem, 2017) 129.6M None
ZSCL (Zheng et al., 2023) 129.6M None
MoE-Adapters (Yu et al., 2024) 59.8M 13.35M
RAIL (Xu et al., 2024) N/A 24.18M / 9.01M
@@
Computation cost. We compare computational costs in terms of trainable parameters and inference overhead, as shown in Table 4. For MoE-Adapter, the gap between training and testing parameters stems from its mixture-of-experts design: all experts are active during training, while only the top-2 are used at inference. RAIL-Primal introduces extra overhead from high-dimensional feature expansion via projection and regression layers. RAIL-Dual further increases memory use by storing image features for all samples across 1,100 classes. These methods all incur linearly growing costs as tasks increase. In contrast, our method reduces training overhead by retaining only high-importance ranks; once merged into the pre-trained weights, the updates add zero memory cost at inference.
@@ 4.4 Ablation Studies
Effects of insertion locations and initial ranks. We study the impact of insertion locations and initial ranks (Fig. 7), consistent with our earlier analyses (Fig. 4). Training restricted to the vision encoder significantly reduces transfer accuracy, indicating degraded pre-trained capabilities. At certain locations, increasing rank improves adaptation but does not always hinder transfer accuracy.
Interestingly, restricting updates to only Attention modules can even surpass our reported state-of-the-art Transfer results (Table 1, Table 3), but at the cost of weaker adaptation. To balance adaptation and retention, we update across all modules by default, achieving an optimal trade-off between learning new tasks and preserving pre-trained knowledge.
Analyses of pruning threshold and dense‐iteration ratio. To understand the trade‐off between learning and forgetting, we vary the maximum threshold $\kappa_{\max}$ and the dense training ratio (Fig. 8). Raising $\kappa_{\max}$ improves Last accuracy by pruning noisy ranks, but overly aggressive pruning limits adaptation. For simplicity, we do not tune the value of $\kappa_{\max}$ for each downstream task. The dense‐iteration ratio controls pruning onset: pruning from the start (ratio = 0) removes ranks before they learn useful knowledge, while delaying too long risks retaining noisy ranks that do not get pruned.
@@ 5 Conclusion
We introduced CoDyRA, a continual learning framework for CLIP that dynamically selects and minimizes LoRA ranks based on their learned importance. Guided by sparsity-promoting regularization, CoDyRA balances plasticity with stability, without storing past data, task information, or relying on assumptions. By keeping updates close to the model’s previous state while preserving essential ranks, our method mitigates forgetting and enhances representation quality. Extensive experiments across diverse continual learning benchmarks demonstrate that CoDyRA achieves state-of-the-art performance, effectively balancing downstream adaptation and knowledge retention.
Future works. CoDyRA updates the model continually by balancing learning and forgetting in the current task. While enabling a flexible and efficient CL process with strong results, future work will further refine the modeling of rank relationships, inter-task dependencies, and forgetting measures.
Reproducibility Statement
Detailed experiment settings and implementation details are provided in Sec. 4.1 and Appendix B.
@@ Appendix A Additional Details of the Proposed Method
We apply proximal gradient descent (Beck & Teboulle, 2009) to handle the non-differentiable $\ell_{1}$ regularization of the importance weights (Ding et al., 2023). At the $t$-th iteration, the update rule for the importance weights of each LoRA becomes:
@@
such that the input to the operator $\mathcal{T}_{\kappa}$ is the elements in $\hat{\mathbf{w}}^{m}_{t}$, $\hat{\mathbf{w}}^{m}_{t}$ is the value of $\mathbf{w}^{m}_{t}$ after the gradient update using only $\mathcal{L}_{\text{sup}}(\Delta_{t})$, and $\kappa=\eta_{t}\cdot\lambda$ is the threshold. The indicator function $\mathbbm{1}(\cdot)$ returns 1 if the condition is met, and 0 otherwise, while $\mathrm{sign}(\cdot)$ denotes the sign (+/)(+/-) of the input. This concludes our derivation of the dynamic update for the importance weights as presented in the main paper Eq. (4).
@@ Appendix B More Experimental Results and Ablations
Further details of the experiment settings. The MTIL setting consists of 1,201 classes drawn from 11 diverse datasets: Aircraft (Maji et al., 2013), Caltech101 (Fei-Fei et al., 2004), CIFAR100 (Krizhevsky et al., 2009), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), Flowers (Nilsback & Zisserman, 2008), Food (Bossard et al., 2014), MNIST (Deng, 2012), OxfordPet (Parkhi et al., 2012), Cars (Krause et al., 2013), and SUN397 (Xiao et al., 2010). In the X-TAIL setting, a total of 10 datasets are used, with CIFAR100 (Fei-Fei et al., 2004) excluded to prevent domain overlap, following the protocol in (Xu et al., 2024). In line with (Xu et al., 2024), we use a 5-shot split for MTIL and a 16-shot split for X-TAIL. For initialization, we follow the standard convention: matrix $\mathbf{A}$ is randomly initialized, matrix $\mathbf{B}$ is set to zero, and the importance weights $\mathbf{w}$ are randomly initialized.
B.1 More Experiment Results
Table 5: Extended results for Table 1, comparing continual fine-tuning with different LoRA settings and the LoRA variant of LAE (Gao et al., 2023).
Method
Aircraft Caltech101 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
CLIP
   Zero-shot 23.5 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 58.5
Transfer
   Zero-shot (Radford et al., 2021) 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   LoRA (rr=4) 74.9 39.4 41.8 67.7 83.3 44.1 87.1 64.2 62.7 62.8
   LoRA (rr=8) 77 38.2 38.4 67.8 83.3 44.6 87.6 64 62.8 62.6
   LoRA (rr=16) 76.4 39.1 37.5 65.9 82.7 41.9 86.9 63.4 62.7 61.8
   LAE-LoRA (Gao et al., 2023) 76.2 39.9 43.8 66.5 83.3 42.7 87.5 63 62.9 62.9
   CoDyRA 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
[…]
Average
   LoRA (rr=4) 23.5 78.4 40.6 47.3 71.4 83.7 50.1 87.5 65.1 64 61.2
   LoRA (rr=8) 23.7 78.1 39.7 46.6 71.3 83.7 50.7 87.8 65.4 64.1 61.1
   LoRA (rr=16) 23.6 78.5 40.5 43.0 70.4 82.8 49.5 87.5 64.9 64.0 60.5
   LAE-LoRA (Gao et al., 2023) 29.2 81.3 50.7 68.8 77.4 83.8 53.2 88.5 66.3 64.2 66.4
   CoDyRA 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
[…]
Last
   LoRA (rr=4) 9.75 79.5 40.7 37.5 65.1 82.6 47.1 84.1 57.5 76.3 58
   LoRA (rr=8) 12.7 78.7 39.6 42.6 62.4 82.3 46.8 83.6 60.2 76.1 58.5
   LoRA (rr=16) 13.7 78.6 40.3 39.4 66.1 82.7 50.9 85.7 60.3 76.1 59.4
   LAE-LoRA (Gao et al., 2023) 25.6 81.5 50.2 73.2 79.5 85 70.4 91.1 72.9 75.7 70.5
   CoDyRA 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
[…]
B.1.1 Comparison to fixed‐rank LoRA baselines.
Table 5 reports results for fixed‐rank LoRA (r = 4, 8, 16) and the LoRA variant of LAE. Three key observations emerge: (1) As LoRA rank increases, Last accuracy improves while Average accuracy declines, aligning with Fig. 4: a relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation. (2) Fixed‐rank LoRA shows a substantial Last‐accuracy gap compared to our adaptive method. This suggests that without dynamic rank selection, low‐relevance ranks (potentially encoding noise) are retained in the model, impairing generalization to unseen domains and accumulating inter‐task interference over time, resulting in severe performance drop. (3) Leveraging fast online LoRA updates, slow EMA aggregation in offline LoRA, and energy-based logit ensembling, LAE substantially outperforms fixed-rank LoRA (r = 16). Nonetheless, unlike our method, fixed-rank LoRA lacks fine-grained control over the contribution of individual ranks within each module.
Table 6: Comparisons on full-shot MTIL setting.
Method
Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
CLIP
   Zero-shot (Radford et al., 2021) 24.3 88.4 68.2 44.6 54.9 71 88.5 59.4 89 64.7 65.2 65.2
Transfer
   Zero-shot (Radford et al., 2021) 74.5 56.9 39.1 51.1 52.6 72.8 60.6 75.1 30.3 55.9 56.9
   LwF (Li & Hoiem, 2017) 56.6 44.6 32.7 39.3 46.6 68.0 46.0 77.4 31.9 60.5 50.4
   iCaRL (Rebuffi et al., 2017a) 77.1 61.0 40.5 45.3 54.4 74.6 47.9 76.7 36.3 58.6 57.2
   LwF-VR (Ding et al., 2022) 73.5 55.6 35.6 41.5 47.0 68.3 53.9 69.3 26.8 51.9 52.3
   WiSE-FT (Wortsman et al., 2022) 86.0 67.4 45.4 50.4 69.1 87.6 61.8 86.8 60.1 66.8 68.1
   ZSCL (Zheng et al., 2023) 55.7 50.9 30.4 41.4 49.3 71.8 36.3 77.5 55.3 53.4 53.2
[…]
Average
   LwF (Li & Hoiem, 2017) 36.3 86.9 72.0 59.0 73.7 60.0 73.6 74.8 80.0 37.3 58.1 64.7
   iCaRL (Rebuffi et al., 2017a) 35.5 89.2 72.2 60.6 68.8 70.0 78.2 62.3 81.8 41.2 62.5 65.7
   LwF-VR (Ding et al., 2022) 29.6 87.7 74.4 59.5 72.4 63.6 77.0 66.7 81.2 43.7 60.7 65.1
   WiSE-FT (Wortsman et al., 2022) 26.7 86.5 64.4 57.1 65.7 58.7 71.1 70.5 75.8 36.9 54.6 60.7
   ZSCL (Zheng et al., 2023) 45.1 92.0 80.1 64.3 79.5 81.6 89.6 75.2 88.9 64.7 68.0 75.4
[…]
Last
   LwF (Li & Hoiem, 2017) 26.3 87.5 71.9 66.6 79.9 66.9 83.8 99.6 92.1 66.1 80.4 74.6
   iCaRL (Rebuffi et al., 2017a) 35.8 93.0 77.0 70.2 83.3 88.5 90.4 86.7 93.2 81.2 81.9 80.1
   LwF-VR (Ding et al., 2022) 20.5 89.8 72.3 67.6 85.5 73.8 85.7 99.6 93.1 73.3 80.9 76.6
   WiSE-FT (Wortsman et al., 2022) 27.2 90.8 68.0 68.9 86.9 74.0 87.6 99.6 92.6 77.8 81.3 77.7
   ZSCL (Zheng et al., 2023) 40.6 92.2 81.3 70.5 94.8 90.5 91.9 98.7 93.9 85.3 80.2 83.6
[…]
B.1.2 Additional results on full-shot MTIL.
To further extend our evaluation, we report full-shot results on the MTIL benchmark (Table 6). For fairness, we adopt the same domain-prediction design used in recent methods. Consistent with Tables 1 and 3, CoDyRA achieves state-of-the-art performance, surpassing all baselines on both Transfer (unseen-domain generalization) and Last (continual downstream learning).
Table 6 also compares against prompt-based approaches (L2P (Wang et al., 2022e), DualPrompt (Wang et al., 2022d), S-Prompt (Wang et al., 2022c)). Although these methods mitigate task forgetting via key–value matching for prompt selection, they neglect the preservation of pre-trained capabilities—evidenced by substantial drops in transfer accuracy—and further underperform recent domain-prediction approaches on last accuracy.
Table 7: Statistical significance of CoDyRA results corresponding to Table 1.
    Aircraft Caltech101 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
CoDyRA Transfer $74.3^{\pm 0.52}$ $36.8^{\pm 0.23}$ $44.2^{\pm 0.56}$ $69.9^{\pm 0.56}$ $83.5^{\pm 0.23}$ $42.8^{\pm 0.18}$ $88.9^{\pm 0.42}$ $64.6^{\pm 0.47}$ $63.4^{\pm 0.56}$ $63.2^{\pm 0.28}$
CoDyRA Average $41.4^{\pm 0.28}$ $81^{\pm 0.38}$ $58.7^{\pm 0.26}$ $77.8^{\pm 0.47}$ $83.4^{\pm 0.39}$ $84.6^{\pm 0.28}$ $64.5^{\pm 0.14}$ $90.4^{\pm 0.4}$ $67.2^{\pm 0.23}$ $64.4^{\pm 0.47}$ $71.3^{\pm 0.18}$
CoDyRA Last $37.7^{\pm 0.42}$ $81.5^{\pm 0.24}$ $65.1^{\pm 0.63}$ $89.9^{\pm 0.55}$ $91.4^{\pm 0.38}$ $85.5^{\pm 0.16}$ $96.8^{\pm 0.08}$ $93.3^{\pm 0.3}$ $77.3^{\pm 0.66}$ $73.5^{\pm 0.21}$ $79.2^{\pm 0.18}$
B.1.3 Training stability.
Table 7 reports the mean and standard deviation of CoDyRA’s results in Table 1, averaged over three independent runs. The consistently low variance across metrics demonstrates that our method is stable to random initialization, highlighting CoDyRA’s robust adaptive rank selection.
B.2 Effects of Additional Adaptive Prediction Head on Visual Representations